# thanks: These two authors contributed equallythanks: These two authors contributed equally

### v5

@@ Abstract
Teaser: Weak ergodicity breaking occurs prior to strong ergodicity breaking in spin and structural glasses.
Introduction
@@
FIG. 1: Basic aging phenomenon in glasses. Data are obtained for the WCA model at $T=0.25$ with $N=8000$ particles, quenched from $T_{\rm eq}=5$ to $T$. (a) $dE/dt_{\rm w}$ as a function of $t_{\rm w}$ at a few different $T$. The solid line indicates $dE/dt_{\rm w}\sim t_{\rm w}^{-1}$, i.e., $E(t_{\rm w})\sim-\ln(t_{\rm w})$. (b) Non-equilibrium decorrelation time $\tau_{\rm neq}$ as a function of $t_{\rm w}$. The solid line represents $\tau_{\rm neq}\sim t_{\rm w}$. (c) $F(t_{\rm w},t_{\rm w}+t)$ for $t_{\rm w}=0.2,2,20,2\times 10^{2},2\times 10^{3},2\times 10^{4},8\times 10^{5}$ (black curves; from left to right), and $F_{w}(t_{\rm w})$ with $w=1/2$ and $w=3/2$. The dashed line indicates the nonergodicity parameter $f$. (c) $F(t_{\rm w},t_{\rm w}+t)$ plotted as a function of $w=t/t_{\rm w}$ for intermediate waiting times $t_{\rm w}=2,20,2\times 10^{2},2\times 10^{3}$. The blue line is $H$ given by Eq. (2) at the given $\hat{x}=T/T_{\rm WEB}$, and the red line is $\tilde{H}=fH$.
@@
Results
Basic aging phenomenon and weak ergodicity breaking
@@ Results
FIG. 2: Aging function $\Pi_{w}(t_{\rm w})$ for the (a,b) E-REM and (c-d) G-REM. In both models, $w=0.5$. Open points are MC simulation data. (a) E-REM results with $N=20$ (points). Data in the plateau regime are fitted by $\Pi(t_{\rm w})=H\left(1+A\,t_{\rm w}^{-\alpha}\right)$ (lines). (b) E-REM data at $T=0.75$. The purple dashed line represents the short-time behavior $\Pi_{w}^{\rm s}(t_{\rm w})=1-Ct_{\rm w}$ with $C\approx 0.214$. The dotted horizontal line represents the BTM plateau $H(w,x)\approx 0.303$. The thermalization time $\tau_{\rm th}=\exp(N/T)$ is marked by crosses. (c) G-REM results with $N=128$ (points). Data at large times are fitted by $\Pi(t_{\rm w})=H\left(1+A\,t_{\rm w}^{-\alpha}\right)-k\ln t_{\rm w}$ (lines). (d) At $T=0.75$, the G-REM data are dominated by logarithmic decay $\Pi(t_{\rm w})\sim-\frac{B}{N}\ln t_{\rm w}$ (lines) in small systems, where $B=0.7$ is a fitting parameter. The horizontal dashed lines represent the GTM plateau $H(w,\hat{x})\approx 0.854$ and the BTM plateau $H(w,x)\approx 0.136$. The fitting parameters $H$ and $\alpha$ for both models are plotted in Fig. 7(a,b).
@@
FIG. 3: Barrier energy distributions in the G-REM and E-REM. (a) A sub-tree formed by local minima and saddle points for the $N=6$ G-REM obtained by the BT algorithm, with corresponding spin configurations and energies ($E$-axis) indicated. (inset) Illustration of two traps: the blue nodes are non-minimum-non-saddle configurations; the barrier energy $\Delta E$ is the energy difference between the local minimum and saddle point. (b) Distributions $p_{\rm G-REM}^{\rm BT}(\Delta E)$ and $p_{\rm E-REM}^{\rm BT}(\Delta E)$ of the G-REM and E-REM by the BT algorithm ($N=20$), compared to $p_{\rm BTM}(\Delta E)$ (Eq. 1) and $p_{\rm GTM}(\Delta E)$ (Eq. 3 with $a=1.12$ and $b=0.96$). (c) Remainder $\ln p(\Delta E)+\Delta E/T_{\rm c}$ for the G-REM (black) and E-REM (red). Open and filled points are obtained by the BT algorithm and TE theory respectively. Lines represent fitting to a Gaussian function. The variance $\sigma^{2}$ and mean $\bar{E}$ obtained by the Gaussian fitting are plotted in (d) as functions of $N$. Linear fitting of the data in (d) gives $a=1.12$ and $b=0.96$ ($\sigma^{2}=aN$ and $\bar{E}=bN$).
Bouchaud’s trap model
@@
FIG. 4: WEB in (a,b) E-REM and (c,d) G-REM. In (a,c), $\Pi_{w}(t_{\rm w})$ is obtained with $N=20$ and $w=0.5$ by MC simulations at the rescaled time $\tilde{t}_{\rm w}=Ct_{\rm w}$ with $C=\frac{w\hat{x}}{1+\hat{x}}$. The solid line is $\Pi^{0}_{w}(\tilde{t}_{\rm w})=1-\tilde{t}_{\rm w}$ (see Eq. 5). The dotted horizontal line is $\Pi_{\rm thr}=0.2$. The time scales $\tilde{\tau}^{0}$ (circle) and $\tilde{\tau}$ (cross) are marked. (b,d) $R=\tilde{\tau}/\tilde{\tau}^{0}$ as a function of $T$ for different NN. (inset) $R$ as a function of NN at $T=1.17\,T_{\rm c}$.
@@
In the REM, it was claimed that $T_{\rm c}$ corresponds to the spin glass transition temperature (see Materials and Methods A for the details of the REM). The exponential form Eq. (1) is consistent with mean-field thermodynamic calculations for the glass models in the so-called the one-step replica symmetry breaking (1-RSB) universality class [1]. Interestingly, Eq. (1) is formally identical to the Boltzmann distribution, $P(E)\sim\exp(-E/T)$, of the microstate energy in an equilibrium system - the only difference is the replacement of $T$ by a constant $T_{\rm c}$. Because $P(\Delta E)$ is a static property of the energy landscape, it should be $T$-independent; then a natural guess of the parameter $T_{\rm c}$ in Eq. (1) is the glass transition temperature, which is the only characteristic temperature in simple models such as the REM.
According to Eq. (1) and the Arrhenius law, $\tau$ follows a power-law distribution, $\psi(\tau)\sim\tau^{-(1+x)}$, where $x=T/T_{\rm c}$ is the reduced temperature. Because $0<x<1$ below $T_{\rm c}$, the average trapping time $\langle\tau\rangle$ diverges in the glass phase. Thus the system takes infinite time to reach equilibrium, leading to long-time aging and WEB. Mathematically, WEB is described by the arcsin law [47]:
@@
FIG. 5: Aging in (a-c) WCA and (d-f) amorphous silica models. (a) $F_{w}(t_{\rm w})$ obtained by MD simulations of the WCA model, with $N=16000$. The data in the intermediate time regime are fitted to $F_{w}(t_{\rm w})=\tilde{H}\left(1+A\,t_{\rm w}^{-\alpha}\right)-k\ln t_{\rm w}$ (solid lines). At the two lowest $T$, $\tilde{H}$ is estimated by fitting the plateau (dashed lines), and $\alpha$ cannot be determined due to the presence of a dip. (b) $F_{w}(t_{\rm w})$ at $T=0.26$ with different NN. Inset is the close-up for the data in $1<t_{\rm w}<10^{3}$, fitted to $F_{w}(t_{\rm w})\sim-k\ln t_{\rm w}$ (lines). The parameter $k$ obtained in this way is plotted in (c) as a function of NN at different $T$. The crossover size $n$ is determined by the intersect of two linear fitting lines (when $T<0.35$). (d-e) Similar plots for amorphous silica (${\rm SiO}_{2}$). In (d,f), $N=31944$; in (e), $T=4000$ K. In all panels, $w=1/2$. The parameters $H$ and $\alpha$ are plotted in Fig. 7(a,b) for both models.
Generalized trap model
To understand why the BTM fails to describe the simulation results, one should naturally re-examine the underlying assumptions. Based on the numerical data obtained for REMs, we have examined three major assumptions as described above: (i) the Arrhenius law $\tau\sim\exp(\Delta E/T)$, (ii) the renewal mechanism $\psi(\tau_{1},\tau_{2})=\psi(\tau_{1})\psi(\tau_{2})$, and (iii) the exponential barrier distribution Eq. (1). We find that the first two assumptions hold, but the third one requires correction (see below). Based on these observations, we propose a GTM, which retains all assumptions of the original BTM, except for the exponential barrier distribution. In the GTM, we add a Gaussian correction term to the barrier distribution,
@@
where $a$ is a model-dependent parameter. The energy scale $\bar{E}=bN$ is assumed to be extensive, where $b$ is also model-dependent. Note that the Gaussian term, which follows the standard central limit theorem, is essentially caused by finite-size effects. Equation (3) will be explicitly examined in the REMs below, but for now, we take it as input and derive the corresponding aging function theoretically (see details in the Supplementary Materials, SM, Sec. S1).
@@
FIG. 6: Static lengths. The following data are plotted: $\xi_{\rm ag}$ for the WCA model (this work), $\xi_{\rm Hessian}$ for the KA-LJ model from [43] ($T$ is rescaled by a factor of $3/4$), $\xi_{\rm PTS}$ for the WCA model from [42], and $\xi_{\rm PTS}$ for HSs from  [50]. The HS $\xi_{\rm PTS}$ is plotted as a function of $(Z^{-1}-Z^{-1}_{\rm K})/(Z_{\rm MCT}^{-1}-Z^{-1}_{\rm K})$. The black line represents the fitting $\xi\sim(T-T_{\rm K})^{-\nu}$, where $T_{\rm K}=0.21$ and $\nu=0.61(1).$ (inset) $\xi$ as a function of $T$.
FIG. 7: Unified phase diagram of activated aging and WEB in glasses. (a) WEB order parameter $H$ and (b) power-law convergence exponent $\alpha$ obtained by simulations of four glass models ($w=t/t_{\rm w}=1/2$). The theoretical lines are Eqs. (2) and (6). (c) In the unified phase diagram, WEB and ergodic phases are separated by the $T/T_{\rm SEB}=1/(1-W)$ line, and the $H$-axis quantifies the order parameter. Each glass model corresponds to one line in the plot with the corresponding $W$. Insets are schematic energy landscapes in ergodic and WEB phases; the arrow indicates hopping to lower energy basins during aging.
Aging in spin glass random energy models
In REMs, we examine both the assumptions and predictions of the GTM. We first focus on the assumptions. As shown in SM Sec. S2, data obtained from numerical simulations are consistent with the Arrhenius law and renewal mechanism. Next we examine the barrier energy distribution, Eq. (3), using two independent approaches. The first one is an exhaustive enumeration of energy barriers using the barrier-tree (BT) algorithm [51, 52, 53, 54] (see Fig. 3a for an example of a sub-tree and Materials and Methods A for algorithm details). The BT method provides the exact distribution $p^{\rm BT}(\Delta E)$, but is restricted to small systems ($N\leq 20$). We also develop a tree-expansion (TE) theory to compute the distribution analytically and approximately, giving $p^{\rm TE}(\Delta E)$ for any $N$ (see SM Sec. S3). The difference between $p^{\rm BT}(\Delta E)$ and $p^{\rm TE}(\Delta E)$ is unnoticeable, as shown in Fig 3(b-d).
@@
An interesting prediction from the GTM theory is that the WEB does not necessarily co-occur with the spin glass transition: for the E-REM, $T_{\rm WEB}=T_{\rm c}$, while for the G-REM, $T_{\rm WEB}\approx 3.6\,T_{\rm c}$. The factor 3.6 in the G-REM is substantially larger than one - thus the GTM scenario ($T_{\rm WEB}\approx 3.6\,T_{\rm c}$) can be unambiguously tested against the original BTM scenario ($T_{\rm WEB}=T_{\rm c}$[22]. Below we directly verify the GTM scenario. In Fig. 4 we plot the MC data of $\Pi_{w}(t_{\rm w})$ for both E-REM and G-REM, with $t_{\rm w}$ rescaled by $C=\frac{w\hat{x}}{1+\hat{x}}$ to collapse the short-time behavior $\Pi_{w}^{\rm 0}(t_{\rm w})=1-Ct_{\rm w}$ at different $T$ (note that $C$ is not a fitting parameter). If WEB occurs, $\Pi_{w}(\tilde{t}_{\rm w})$ should not vanish in the $\tilde{t}_{\rm w}\to\infty$ limit, where $\tilde{t}_{\rm w}\equiv C\,{t}_{\rm w}$ is the reduced waiting time. We define a time scale $\tilde{\tau}$ such that $\Pi_{w}(\tilde{t}_{\rm w}=\tilde{\tau})=\Pi_{\rm thr}$ and choose a threshold $\Pi_{\rm thr}=0.2$. At sufficiently high $T$ (in the ergodic phase), $\tilde{\tau}=\tilde{\tau}^{0}$ is a constant, where $\tilde{\tau}^{0}$ is defined by $\Pi_{w}^{0}(\tilde{t}_{\rm w}=\tilde{\tau}^{0})=1-\tilde{\tau}^{0}=\Pi_{\rm thr}$. As shown in Fig. 4, with an increasing $T$, the ratio $R=\tilde{\tau}/\tilde{\tau}^{0}$ approaches one around $T_{\rm c}$ in E-REM and around $T_{\rm WEB}\approx 3.6\,T_{\rm c}$ in G-REM. According to the GTM scenario, we expect $R\to\infty$ below $T_{\rm WEB}$ in the thermodynamic limit, because $\tilde{\tau}$ should diverge in the WEB phase. This expectation is confirmed by finite-size analyses at a fixed $T=1.17\,T_{\rm c}$ slightly above $T_{\rm c}$ in both models (see Fig. 4d-inset). In G-REM, $R$ increases with $N$ at this temperature; in contrast, in E-REM, $R\approx 1$ is nearly invariant. This is direct evidence of WEB above $T_{\rm c}$ in the G-REM.
Table 1: Characteristic temperatures. The spin glass transition temperature $T_{\rm c}$ is obtained by thermodynamic calculations for G-REM [33] and E-REM [55]. The divergence temperature $T_{\rm 0}$ is estimated from the VFT fitting of the $\alpha$-relaxation time $\tau_{\rm eq}\sim\exp\left(\frac{C}{T-T_{\rm 0}}\right)$ for the WCA model [56]. The MCT temperature $T_{\rm MCT}$ is estimated from the MCT fitting, $\tau_{\rm eq}\sim(T-T_{\rm MCT})^{-\gamma}$ for the WCA model [57], and $D\sim(T-T_{\rm MCT})^{\gamma}$ of the diffusion constant $D$ for amorphous $\rm{SiO}_{2}$ [58]. The onset temperature $T_{\rm onset}$ in the WCA model corresponds to the crossover temperature where the Stokes-Einstein relation, $\tau_{\rm eq}\sim D^{-1}$, breaks down [59]. Using four different criteria, Ref. [60] reports $T_{\rm onset}=0.6-0.7$ for the WCA model. Data without references are obtained by the current study (see the main text and SM Sec. S4 for details).
@@
Aging in structural glasses
To further examine the universality of the GTM theory, we simulate two structural glass formers: the WCA model as an example of simple molecular systems, and the amorphous silica ($\rm{SiO}_{2}$) model as an example of network glasses (see Materials and Methods B and C for details). Because efficient sampling of energy barriers in structural glasses remains technically challenging [40, 61, 62], we will only compare the simulated aging function $\Pi_{w}(t_{\rm w})$ to theoretical predictions (see the data in Fig. 5 with a fixed $w=1/2$). Note that $\Pi_{w}(t_{\rm w})$ is related to $F_{w}(t_{\rm w})$ via $F_{w}(t_{\rm w})=f\Pi_{w}(t_{\rm w})$, where $f$ is the nonergodicity parameter defined by the large-time plateau value of the equilibrium incoherent scattering function $f=\lim_{t\to\infty}F_{\rm eq}(t)$. The nonergodicity parameter $f$ is analogous to the Edwards-Anderson order parameter $q_{\rm EA}$ in spin glasses, both of which are determined by the typical size of the glass basins.
We first extract $H$ and $\alpha$ from the aging data, and compare them to theoretical predictions. As shown in Fig. 5(a,d), the high-$T$ simulation results can be fitted to the form, $F_{w}(t_{\rm w})=\tilde{H}\left(1+A\,t_{\rm w}^{-\alpha}\right)-k\ln t_{\rm w}$, from which $\tilde{H}$ and $\alpha$ are obtained. Note that this expression only works in the intermediate-$t_{\rm w}$ regime (see Eq. 5), and thus the small-$t_{\rm w}$ (microscopic dynamics) and large-$t_{\rm w}$ (thermalization) data points should be excluded from the fitting. At lower $T$, a dip in $F_{w}(t_{\rm w})$, which might be attributed to the so-called boson peak [63], appears after the short-time processes, making the above fitting form inapplicable. Nevertheless, at these temperatures, $\tilde{H}$ can still be estimated since the plateau is well-defined. Then the relation $H=\tilde{H}/f$ is used to estimate $H$ (see SM Sec. S4 for the data of $f$). The values of $H$ and $\alpha$ are plotted in Fig. 7(a,b), as functions of $\hat{x}=T/T_{\rm WEB}$ with a fixed $w=1/2$. With a proper choice of $T_{\rm WEB}$, these data collapse with the GTM theoretical curves Eq. (2) and (6). Note that $T_{\rm WEB}$ is obtained from fitting the data to the arcsin law Eq. (2). A further consistent check is illustrated in Fig. 1d for the WCA model: using the same $T_{\rm WEB}$ and $f$, the theoretical curve $\tilde{H}(w,\hat{x}=0.36)=fH(w,\hat{x}=0.36)$, with $H$ given by Eq. (2), agrees with the simulation data $F^{\rm ag}(w)$. Together, Fig. 1d and 7(a,b) suggest that the theory works for different $T$ and $w=t/t_{\rm w}$.

### v6

# Universal activated aging and weak ergodicity breaking in spin and structural glasses

@@ Abstract
I Introduction
@@
Figure 1: Basic aging phenomenon in glasses. Data are obtained for the WCA model at $T=0.25$ with $N=8000$ particles, quenched from $T_{\rm eq}=5$ to $T$. (a) $dE/dt_{\rm w}$ as a function of $t_{\rm w}$ at a few different $T$. The solid line indicates $dE/dt_{\rm w}\sim t_{\rm w}^{-1}$, i.e., $E(t_{\rm w})\sim-\ln(t_{\rm w})$. (b) Non-equilibrium decorrelation time $\tau_{\rm neq}$ as a function of $t_{\rm w}$. The solid line represents $\tau_{\rm neq}\sim t_{\rm w}$. (c) $F(t_{\rm w},t_{\rm w}+t)$ for $t_{\rm w}=0.2,2,20,2\times 10^{2},2\times 10^{3},2\times 10^{4},8\times 10^{5}$ (black curves; from left to right), and $F_{w}(t_{\rm w})$ with $w=1/2$ and $w=3/2$. The dashed line indicates the nonergodicity parameter $f$. (c) $F(t_{\rm w},t_{\rm w}+t)$ plotted as a function of $w=t/t_{\rm w}$ for intermediate waiting times $t_{\rm w}=2,20,2\times 10^{2},2\times 10^{3}$. The blue line is $H$ given by Eq. (2) at the given $\hat{x}=T/T_{\rm WEB}$, and the red line is $\tilde{H}=fH$.
@@
II Basic aging phenomenon and weak ergodicity breaking
@@
Figure 2: Barrier energy distributions in the G-REM and E-REM. (a) A sub-tree formed by local minima and saddle points for the $N=6$ G-REM obtained by the BT algorithm, with corresponding spin configurations and energies ($E$-axis) indicated. (inset) Illustration of two traps: the blue nodes are non-minimum-non-saddle configurations; the barrier energy $\Delta E$ is the energy difference between the local minimum and saddle point. (b) Distributions $p_{\rm G-REM}^{\rm BT}(\Delta E)$ and $p_{\rm E-REM}^{\rm BT}(\Delta E)$ of the G-REM and E-REM by the BT algorithm ($N=20$), compared to $p_{\rm BTM}(\Delta E)$ (Eq. 1) and $p_{\rm GTM}(\Delta E)$ (Eq. 3 with $a=1.12$ and $b=0.96$). (c) Remainder $\ln p(\Delta E)+\Delta E/T_{\rm c}$ for the G-REM (black) and E-REM (red). Open and filled points are obtained by the BT algorithm and TE theory respectively. Lines represent fitting to a Gaussian function. The variance $\sigma^{2}$ and mean $\bar{E}$ obtained by the Gaussian fitting are plotted in (d) as functions of $N$. Linear fitting of the data in (d) gives $a=1.12$ and $b=0.96$ ($\sigma^{2}=aN$ and $\bar{E}=bN$).
@@
Figure 3: Aging function $\Pi_{w}(t_{\rm w})$ for the (a,b) E-REM and (c-d) G-REM. In both models, $w=0.5$. Open points are MC simulation data. (a) E-REM results with $N=20$ (points). Data in the plateau regime are fitted by $\Pi(t_{\rm w})=H\left(1+A\,t_{\rm w}^{-\alpha}\right)$ (lines). (b) E-REM data at $T=0.75$. The purple dashed line represents the short-time behavior $\Pi_{w}^{\rm s}(t_{\rm w})=1-Ct_{\rm w}$ with $C\approx 0.214$. The dotted horizontal line represents the BTM plateau $H(w,x)\approx 0.303$. The thermalization time $\tau_{\rm th}=\exp(N/T)$ is marked by crosses. (c) G-REM results with $N=128$ (points). Data at large times are fitted by $\Pi(t_{\rm w})=H\left(1+A\,t_{\rm w}^{-\alpha}\right)-k\ln t_{\rm w}$ (lines). (d) At $T=0.75$, the G-REM data are dominated by logarithmic decay $\Pi(t_{\rm w})\sim-\frac{B}{N}\ln t_{\rm w}$ (lines) in small systems, where $B=0.7$ is a fitting parameter. The horizontal dashed lines represent the GTM plateau $H(w,\hat{x})\approx 0.854$ and the BTM plateau $H(w,x)\approx 0.136$. The fitting parameters $H$ and $\alpha$ for both models are plotted in Fig. 7(a,b).
III Bouchaud’s trap model
@@
Figure 4: WEB in (a,b) E-REM and (c,d) G-REM. In (a,c), $\Pi_{w}(t_{\rm w})$ is obtained with $N=20$ and $w=0.5$ by MC simulations at the rescaled time $\tilde{t}_{\rm w}=Ct_{\rm w}$ with $C=\frac{w\hat{x}}{1+\hat{x}}$. The solid line is $\Pi^{0}_{w}(\tilde{t}_{\rm w})=1-\tilde{t}_{\rm w}$ (see Eq. 5). The dotted horizontal line is $\Pi_{\rm thr}=0.2$. The time scales $\tilde{\tau}^{0}$ (circle) and $\tilde{\tau}$ (cross) are marked. (b,d) $R=\tilde{\tau}/\tilde{\tau}^{0}$ as a function of $T$ for different NN. (inset) $R$ as a function of NN at $T=1.17\,T_{\rm c}$.
@@
In the REM, it was claimed that $T_{\rm c}$ corresponds to the spin glass transition temperature (see Appendix A for the details of the REM). The exponential form Eq. (1) is consistent with mean-field thermodynamic calculations for the glass models in the so-called the one-step replica symmetry breaking (1-RSB) universality class [1]. Interestingly, Eq. (1) is formally identical to the Boltzmann distribution, $P(E)\sim\exp(-E/T)$, of the microstate energy in an equilibrium system - the only difference is the replacement of $T$ by a constant $T_{\rm c}$. Because $P(\Delta E)$ is a static property of the energy landscape, it should be $T$-independent; then a natural guess of the parameter $T_{\rm c}$ in Eq. (1) is the glass transition temperature, which is the only characteristic temperature in simple models such as the REM.
According to Eq. (1) and the Arrhenius law, $\tau$ follows a power-law distribution, $\psi(\tau)\sim\tau^{-(1+x)}$, where $x=T/T_{\rm c}$ is the reduced temperature. Because $0<x<1$ below $T_{\rm c}$, the average trapping time $\langle\tau\rangle$ diverges in the glass phase. Thus the system takes infinite time to reach equilibrium, leading to long-time aging and WEB. Mathematically, WEB is described by the arcsin law [47] (see Appendix B):
@@
Figure 5: Aging in (a-c) WCA and (d-f) amorphous silica models. (a) $F_{w}(t_{\rm w})$ obtained by MD simulations of the WCA model, with $N=16000$. The data in the intermediate time regime are fitted to $F_{w}(t_{\rm w})=\tilde{H}\left(1+A\,t_{\rm w}^{-\alpha}\right)-k\ln t_{\rm w}$ (solid lines). At the two lowest $T$, $\tilde{H}$ is estimated by fitting the plateau (dashed lines), and $\alpha$ cannot be determined due to the presence of a dip. (b) $F_{w}(t_{\rm w})$ at $T=0.26$ with different NN. Inset is the close-up for the data in $1<t_{\rm w}<10^{3}$, fitted to $F_{w}(t_{\rm w})\sim-k\ln t_{\rm w}$ (lines). The parameter $k$ obtained in this way is plotted in (c) as a function of NN at different $T$. The crossover size $n$ is determined by the intersect of two linear fitting lines (when $T<0.35$). (d-e) Similar plots for amorphous silica (${\rm SiO}_{2}$). In (d,f), $N=31944$; in (e), $T=4000$ K. In all panels, $w=1/2$. The parameters $H$ and $\alpha$ are plotted in Fig. 7(a,b) for both models.
IV Generalized trap model
To understand why the BTM fails to describe the simulation results, one should naturally re-examine the underlying assumptions. Based on the numerical data obtained for REMs, we have examined three major assumptions as described above: (i) the Arrhenius law $\tau\sim\exp(\Delta E/T)$, (ii) the renewal mechanism $\psi(\tau_{1},\tau_{2})=\psi(\tau_{1})\psi(\tau_{2})$, and (iii) the exponential barrier distribution Eq. (1). We find that the first two assumptions hold (see Appendix A), but the third one requires correction (see below). Based on these observations, we propose a GTM, which retains all assumptions of the original BTM, except for the exponential barrier distribution. In the GTM, we add a Gaussian correction term to the barrier distribution,
@@ IV Generalized trap model
where $a$ is a model-dependent parameter. The energy scale $\bar{E}=bN$ is assumed to be extensive, where $b$ is also model-dependent. Note that the Gaussian term, which follows the standard central limit theorem, is essentially caused by finite-size effects. Equation (3) will be explicitly examined in the REMs below, but for now, we take it as input and derive the corresponding aging function theoretically (see details in the Supplemental Material, SM, Sec. S1).
@@
Figure 6: Static lengths. The following data are plotted: $\xi_{\rm ag}$ for the WCA model (this work), $\xi_{\rm Hessian}$ for the KA-LJ model from [43] ($T$ is rescaled by a factor of $3/4$), $\xi_{\rm PTS}$ for the WCA model from [42], and $\xi_{\rm PTS}$ for HSs from  [50]. The HS $\xi_{\rm PTS}$ is plotted as a function of $(Z^{-1}-Z^{-1}_{\rm K})/(Z_{\rm MCT}^{-1}-Z^{-1}_{\rm K})$. The black line represents the fitting $\xi\sim(T-T_{\rm K})^{-\nu}$, where $T_{\rm K}=0.21$ and $\nu=0.61(1).$ (inset) $\xi$ as a function of $T$.
Figure 7: Unified phase diagram of activated aging and WEB in glasses. (a) WEB order parameter $H$ and (b) power-law convergence exponent $\alpha$ obtained by simulations of four glass models ($w=t/t_{\rm w}=1/2$). The theoretical lines are Eqs. (2) and (6). (c) In the unified phase diagram, WEB and ergodic phases are separated by the $T/T_{\rm SEB}=1/(1-W)$ line, and the $H$-axis quantifies the order parameter. Each glass model corresponds to one line in the plot with the corresponding $W$. Insets are schematic energy landscapes in ergodic and WEB phases; the arrow indicates hopping to lower energy basins during aging.
V Aging in spin glass random energy models
In REMs, we examine both the assumptions and predictions of the GTM. We first focus on the assumptions. As shown in Appendix A, data obtained from numerical simulations are consistent with the Arrhenius law and renewal mechanism. Next we examine the barrier energy distribution, Eq. (3), using two independent approaches. The first one is an exhaustive enumeration of energy barriers using the barrier-tree (BT) algorithm [51, 52, 53, 54] (see Fig. 2a for an example of a sub-tree and Appendix A for algorithm details). The BT method provides the exact distribution $p^{\rm BT}(\Delta E)$, but is restricted to small systems ($N\leq 20$). We also develop a tree-expansion (TE) theory to compute the distribution analytically and approximately, giving $p^{\rm TE}(\Delta E)$ for any $N$ (see SM Sec. S2). The difference between $p^{\rm BT}(\Delta E)$ and $p^{\rm TE}(\Delta E)$ is unnoticeable, as shown in Fig 2(b-d).
@@ V Aging in spin glass random energy models
An interesting prediction from the GTM theory is that the WEB does not necessarily co-occur with the spin glass transition: for the E-REM, $T_{\rm WEB}=T_{\rm c}$, while for the G-REM, $T_{\rm WEB}\approx 3.6\,T_{\rm c}$. The factor 3.6 in the G-REM is significantly larger than one - thus the GTM scenario ($T_{\rm WEB}\approx 3.6\,T_{\rm c}$) can be unambiguously tested against the original BTM scenario ($T_{\rm WEB}=T_{\rm c}$[22]. Below we directly verify the GTM scenario. In Fig. 4 we plot the MC data of $\Pi_{w}(t_{\rm w})$ for both E-REM and G-REM, with $t_{\rm w}$ rescaled by $C=\frac{w\hat{x}}{1+\hat{x}}$ to collapse the short-time behavior $\Pi_{w}^{\rm 0}(t_{\rm w})=1-Ct_{\rm w}$ at different $T$ (note that $C$ is not a fitting parameter). If WEB occurs, $\Pi_{w}(\tilde{t}_{\rm w})$ should not vanish in the $\tilde{t}_{\rm w}\to\infty$ limit, where $\tilde{t}_{\rm w}\equiv C\,{t}_{\rm w}$ is the reduced waiting time. We define a time scale $\tilde{\tau}$ such that $\Pi_{w}(\tilde{t}_{\rm w}=\tilde{\tau})=\Pi_{\rm thr}$ and choose a threshold $\Pi_{\rm thr}=0.2$. At sufficiently high $T$ (in the ergodic phase), $\tilde{\tau}=\tilde{\tau}^{0}$ is a constant, where $\tilde{\tau}^{0}$ is defined by $\Pi_{w}^{0}(\tilde{t}_{\rm w}=\tilde{\tau}^{0})=1-\tilde{\tau}^{0}=\Pi_{\rm thr}$. As shown in Fig. 4, with an increasing $T$, the ratio $R=\tilde{\tau}/\tilde{\tau}^{0}$ approaches one around $T_{\rm c}$ in E-REM and around $T_{\rm WEB}\approx 3.6\,T_{\rm c}$ in G-REM. According to the GTM scenario, we expect $R\to\infty$ below $T_{\rm WEB}$ in the thermodynamic limit, because $\tilde{\tau}$ should diverge in the WEB phase. This expectation is confirmed by finite-size analyses at a fixed $T=1.17\,T_{\rm c}$ slightly above $T_{\rm c}$ in both models (see Fig. 4d-inset). In G-REM, $R$ increases with $N$ at this temperature; in contrast, in E-REM, $R\approx 1$ is nearly invariant. This is direct evidence of WEB above $T_{\rm c}$ in the G-REM.
Table 1: Characteristic temperatures. The spin glass transition temperature $T_{\rm c}$ is obtained by thermodynamic calculations for G-REM [33] and E-REM [55]. The divergence temperature $T_{\rm 0}$ is estimated from the VFT fitting of the $\alpha$-relaxation time $\tau_{\rm eq}\sim\exp\left(\frac{C}{T-T_{\rm 0}}\right)$ for the WCA model [56]. The MCT temperature $T_{\rm MCT}$ is estimated from the MCT fitting, $\tau_{\rm eq}\sim(T-T_{\rm MCT})^{-\gamma}$ for the WCA model [57], and $D\sim(T-T_{\rm MCT})^{\gamma}$ of the diffusion constant $D$ for amorphous $\rm{SiO}_{2}$ [58]. The onset temperature $T_{\rm onset}$ in the WCA model corresponds to the crossover temperature where the Stokes-Einstein relation, $\tau_{\rm eq}\sim D^{-1}$, breaks down [59]. Using four different criteria, Ref. [60] reports $T_{\rm onset}=0.6-0.7$ for the WCA model. Data without references are obtained by the current study (see the main text and Appendix for details).
@@
VI Aging in structural glasses
To further examine the universality of the GTM theory, we simulate two structural glass formers: the WCA model as an example of simple molecular systems, and the amorphous silica ($\rm{SiO}_{2}$) model as an example of network glasses (see Appendixes C and D for details). Because efficient sampling of energy barriers in structural glasses remains technically challenging [40, 61, 62], we will only compare the simulated aging function $\Pi_{w}(t_{\rm w})$ to theoretical predictions (see the data in Fig. 5 with a fixed $w=1/2$). Note that $\Pi_{w}(t_{\rm w})$ is related to $F_{w}(t_{\rm w})$ via $F_{w}(t_{\rm w})=f\Pi_{w}(t_{\rm w})$, where $f$ is the nonergodicity parameter defined by the large-time plateau value of the equilibrium incoherent scattering function $f=\lim_{t\to\infty}F_{\rm eq}(t)$. The nonergodicity parameter $f$ is analogous to the Edwards-Anderson order parameter $q_{\rm EA}$ in spin glasses, both of which are determined by the typical size of the glass basins.
We first extract $H$ and $\alpha$ from the aging data, and compare them to theoretical predictions. As shown in Fig. 5(a,d), the high-$T$ simulation results can be fitted to the form, $F_{w}(t_{\rm w})=\tilde{H}\left(1+A\,t_{\rm w}^{-\alpha}\right)-k\ln t_{\rm w}$, from which $\tilde{H}$ and $\alpha$ are obtained. Note that this expression only works in the intermediate-$t_{\rm w}$ regime (see Eq. 5), and thus the small-$t_{\rm w}$ (microscopic dynamics) and large-$t_{\rm w}$ (thermalization) data points should be excluded from the fitting. At lower $T$, a dip in $F_{w}(t_{\rm w})$, which might be attributed to the so-called boson peak [63], appears after the short-time processes, making the above fitting form inapplicable. Nevertheless, at these temperatures, $\tilde{H}$ can still be estimated since the plateau is well-defined. Then the relation $H=\tilde{H}/f$ is used to estimate $H$ (see Appendix E for the data of $f$). The values of $H$ and $\alpha$ are plotted in Fig. 7(a,b), as functions of $\hat{x}=T/T_{\rm WEB}$ with a fixed $w=1/2$. With a proper choice of $T_{\rm WEB}$, these data collapse with the GTM theoretical curves Eq. (2) and (6). Note that $T_{\rm WEB}$ is obtained from fitting the data to the arcsin law Eq. (2). A further consistent check is illustrated in Fig. 1d for the WCA model: using the same $T_{\rm WEB}$ and $f$, the theoretical curve $\tilde{H}(w,\hat{x}=0.36)=fH(w,\hat{x}=0.36)$, with $H$ given by Eq. (2), agrees with the simulation data $F^{\rm ag}(w)$. Together, Fig. 1d and 7(a,b) suggest that the theory works for different $T$ and $w=t/t_{\rm w}$.